Type of Donut

Number purchased

Probability

Cumulative Probability

Chocolate Cream filled

5.00

Strawberry

2.00

Kale frosted

9.00

Blueberry

3.00

Cream filled surprise

5.00

Based on the above chart: What is the probability of selling a Blueberry donut?
Answer #1

probability of selling a Blueberry donut

P(blueberry donut) = (no. of blueberry donuts sold)/(total donuts sold)

= 3 / (5+2+9+3+5)

= 3/24

= 0.125

ANSWER : P(blueberry donut) = 0.125

(please upvote)
Answer #2

Calculation:

  1. Total donuts sold = 5 (Chocolate Cream) + 2 (Strawberry) + 9 (Kale frosted) +                                    3 (Blueberry) + 5 (Cream filled surprise) = 24 donuts.


  2. Probability for Blueberry = (Number of Blueberry donuts / Total donuts) ×                                                      100


  3.                                             = (3 / 24) × 100 = 12.5%.


So, there's a 12.5% chance a customer will buy a Blueberry donut.


source: Self
answered by: Harshwardhan kunal